# How individual vs shared coordination governs the degree of correlation in rotational vs residence times in a high-viscosity lithium electrolyte

### v2

@@ Abstract
Commercially used carbonate-based electrolytes in lithium-ion batteries are susceptible to many challenges, including flammability, volatility, and lower thermal stability. Solvated ionic liquids of LiTFSI salt (lithium bis(trifluoromethylsulfonyl)-amide) and glyme-based solvents are potential alternative candidates for commonly used electrolytes. We perform classical molecular dynamics (MD) simulations study the effect of concentration and temperature on the translational and rotational dynamics. The radial distribution function shows stronger coordination of Li+ ions with tetraglyme(G4), as shown in earlier studies, and forms a stable [Li(G4)]+ cation complex. The self-diffusion coefficients are lower than the values experimentally observed but show better improvement over other classical force fields. An increase in the salt concentrations leads to a higher viscosity of the system and reduces the overall ionic mobility of Li+ ions. Diluting the system with a larger number of glyme molecules leads to shorter rotational relaxation times for both TFSI and tetraglyme. Ion-residence times show that Li+ ions form stable and long-lasting complexes with G4 molecules than TFSI anions. The residence time of [Li(G4)]+ complex increases at higher salt concentrations due to the availability of fewer G4 molecules to coordinate with a Li+ ion. G4 is also seen to form polydentate complexes with Li+ without a shared coordination, allowing rotation without breaking coordination, unlike TFSI, which requires coordination disruption for rotation. This distinction explains the poor correlation between rotation and residence time for G4 and the strong correlation for TFSI.
@@ III.2 Radius of gyration
1:4 | TFSI | 0.2564 | 0.0006 | 0.2568 | 0.0006 |
G4 | 0.3808 | 0.0016 | 0.3886 | 0.0018 | |
2:1 | TFSI | 0.2540 | 0.0002 | 0.2550 | 0.0003 |
G4 | 0.3628 | 0.0007 | 0.3672 | 0.0034 |
Since both the TFSI anion and G4 are long molecular structures, their coiling leads to a significant variation in the radii of gyration (). The of these species is an important factor that is sensitive to the temperature and concentration of electrolyte mixture. We used MD trajectories to calculate for TFSI anions and G4 molecules for each ratio and at 300 K and 500 K (Fig. S1 and S2 in supplementary information). As shown in Table 1, the of TFSI and G4 in the most diluted system (1:4) at 300 K is, 0.256 nm and 0.381 nm, respectively. At a higher concentration, 2:1, the for both TFSI and G4 decreases to 0.254 nm and 0.363 nm, suggesting a more coiled configuration, particularly for G4, which shrinks by 5 %. The for both TFSI and G4 decrease with concentration of LiTFSI. This indicates that at higher salt concentrations, most G4 molecules coordinate individually with the Li+ ions, thus increasing their overall compactness. At lower salt concentrations, there are more free G4 molecules available to share the coordination leading to higher . As expected at higher temperature, for TFSI and G4 increase from 0.2564 nm to 0.2568 nm and from 0.3808 nm to 0.3886 nm respectively, when increasing the temperature from 300 K to 500 K. Even these slight changes in is an important factor which can lead to an interplay in translational and rotational dynamics and ion-residence times in these mixtures (discussed vide infra).
@@ III.5 Ion-residence times
for Li…O(G4) and Li…O(TFSI) interactions is calculated within the first solvation shell of lithium. The radius of the solvation shell is taken from the first minima of the corresponding RDFs. Fig. S6 in supplementary information shows the decay of IEAF for Li-O, for both TFSI and G4 at 300 K and 500 K. The decay is too sluggish for Li…O(TFSI) and Li…O(G4) in the systems with higher salt concentration. For all these systems is calculated numerically (for well-decayed) and analytically (for sluggish-decayed) tabulated in Table S3 on the supplementary information. The value of is 6 s for Li…TFSI and 30 s for Li…G4 for 2:1 mixture at 300 K, which decreases to 1 ns and 882 ns, respectively, as the system is diluted with more G4 (1:4). Dong et al.Dong and Bedrov (2018) calculated the residence time of Li…O(TFSI) at various temperatures and showed that, for the 1:1 system, the residence time for Li…O(TFSI) is 7 ns at 303 K and decreases with increasing the concentration of free G4 molecules and at higher temperatures. In our study, for equimolar system, the residence time of Li…O(TFSI) is 445 ns at 300 K and decreased to 1 ns at 500 K, as shown in Table S3 in the supplementary information. Similarly, the residence time of Li…O(G4) is 34 s at 300 K and decreases by three orders to 46 ns at 500 K. In a similar study for an equimolar LITFSI/G4, Shinoda et al.Shinoda et al. (2018) reported that the residence time of Li…TFSI and Li…G4 at 503 K was 2.5 ns and 87 ns, respectively. More dilution leads to an increase in the number of free G4 molecules in the system, enabling their rapid exchange with the Li+ ion coordination shell, subsequently decreasing the residence time.

### v3

@@ Abstract
Commercially used carbonate-based electrolytes in lithium-ion batteries are susceptible to many challenges, including flammability, volatility, and lower thermal stability compared to ionic-liquid based electrolytes. Ionic liquids of LiTFSI salt (lithium bis(trifluoromethylsulfonyl)-amide) and glyme-based solvents are potential alternative candidates for commonly used electrolytes. We perform classical molecular dynamics (MD) simulations to study the effect of concentration and temperature on the translational and rotational dynamics. The radial distribution function shows stronger coordination of Li+ ions with tetraglyme(G4), as shown in earlier studies, and forms a stable [Li(G4)]+ cation complex. The self-diffusion coefficients are lower than the values experimentally observed but shows improvement over other classical force fields, without charge scaling. An increase in the salt concentrations leads to a higher viscosity of the system and reduces the overall ionic mobility of Li+ ions. Diluting the system with a larger number of glyme molecules leads to shorter rotational relaxation times for both TFSI and tetraglyme. Ion-residence times show that Li+ ions form stable and long-lasting complexes with G4 molecules than TFSI anions. The residence time of [Li(G4)]+ complex increases at higher salt concentrations due to the availability of fewer G4 molecules to coordinate with a Li+ ion. G4 is also seen to form polydentate complexes with Li+ ions without a shared coordination, allowing rotation without breaking coordination, unlike TFSI, which requires coordination disruption for rotation. This distinction explains the poor correlation between rotation and residence time for G4 and the strong correlation for TFSI.
@@ III.2 Radius of gyration
1:4 | TFSI | 0.256 | 0.0006 | 0.257 | 0.0006 |
G4 | 0.381 | 0.0016 | 0.389 | 0.0018 | |
2:1 | TFSI | 0.254 | 0.0002 | 0.255 | 0.0003 |
G4 | 0.363 | 0.0007 | 0.367 | 0.0034 |
Since both the TFSI anion and G4 are long molecular structures, their coiling leads to a significant variation in the radii of gyration (). The of these species is an important factor that is sensitive to the temperature and concentration of electrolyte mixture. We used MD trajectories to calculate for TFSI anions and G4 molecules for each ratio and at 300 K and 500 K (Fig. S1 and S2 in supplementary information). As shown in Table 1, the of TFSI and G4 in the most diluted system (1:4) at 300 K is, 0.256 nm and 0.381 nm, respectively. At a higher concentration, 2:1, the for both TFSI and G4 decreases to 0.254 nm and 0.363 nm, suggesting a more coiled configuration, particularly for G4, which shrinks by 5 %. The for both TFSI and G4 decrease with concentration of LiTFSI. This indicates that at higher salt concentrations, most G4 molecules coordinate individually with the Li+ ions, thus increasing their overall compactness. At lower salt concentrations, there are more free G4 molecules available to share the coordination leading to higher . As expected at higher temperature, for TFSI and G4 increase from 0.256 nm to 0.257 nm and from 0.381 nm to 0.387 nm respectively, when increasing the temperature from 300 K to 500 K. Even these slight changes in is an important factor which can lead to an interplay in translational and rotational dynamics and ion-residence times in these mixtures (discussed vide infra).
@@ III.5 Ion-residence times
for Li…O(G4) and Li…O(TFSI) interactions is calculated within the first solvation shell of lithium. The radius of the solvation shell is taken from the first minima of the corresponding RDFs. Fig. S6 in supplementary information shows the decay of IEAF for Li-O, for both TFSI and G4 at 300 K and 500 K. The decay is too sluggish for Li…O(TFSI) and Li…O(G4) in the systems with higher salt concentration. For all these systems is calculated numerically (for well-decayed) and analytically (for sluggish-decayed) tabulated in Table S3 on the supplementary information. The value of is 6 s for Li…TFSI and 30 s for Li…G4 for 2:1 mixture at 300 K, which decreases to 1 ns and 882 ns, respectively, as the system is diluted with more G4 (1:4). Dong et al.Dong and Bedrov (2018) calculated the residence time of Li…O(TFSI) at various temperatures and showed that, for the 1:1 system, the residence time for Li…O(TFSI) is 7 ns at 303 K and decreases with increasing the concentration of free G4 molecules and at higher temperatures. In our study, for equimolar system, the residence time of Li…O(TFSI) is 445 ns at 300 K and decreased to 1 ns at 500 K, as shown in Table S3 in the supplementary information. Similarly, the residence time of Li…O(G4) is 34 s at 300 K and decreases by three orders to 46 ns at 500 K. The discrepancy in the residence time between our study and Dong’s group could be due to the differences in force fields and the treatment of partial charges. They used polarizable force field which capture the dynamic molecular charge redistribution, with the changing environment, which can enhance ion mobility and affect the residence time. In a similar study for an equimolar LITFSI/G4, Shinoda et al.Shinoda et al. (2018) reported that the residence time of Li…TFSI and Li…G4 at 503 K was 2.5 ns and 87 ns, respectively, which is closure to the values we observed. The slight differences may be due to differences in how partial charges were assigned. More dilution leads to an increase in the number of free G4 molecules in the system, enabling their rapid exchange with the Li+ ion coordination shell, subsequently decreasing the residence time.